OpenStudy (anonymous):

30% of 60 is what number?

OpenStudy (jdoe0001):

hint: \(\bf {\color{brown}{ a}}\%\ of\ b \implies \cfrac{{\color{brown}{ a}}}{100}\cdot b\implies {\color{brown}{ a}}\div 100\cdot b\)

OpenStudy (anonymous):

18.. 30/100 * 60 = 18

OpenStudy (anonymous):

Oh, that's what I got out of it. Sorry, don't listen to me. @MrNood will handle this.